Find the equation of the plane through the points (1,-2,4) ,(3,1,-2) and (2,-1,6)


1
Expert's answer
2021-08-18T07:40:28-0400

"\text{The equation of the plane is given by}
\\(\overrightarrow{r} - \overrightarrow{a}).\overrightarrow{n}=0
\\\overrightarrow{r}.\overrightarrow{n}=\overrightarrow{a}.\overrightarrow{n}
\\\text{where $\overrightarrow{a}=(1,-2,4), \overrightarrow{r} = (x,y,z)$}\\\text{Next, we compute $\overrightarrow{n}$, to this we compute the following vectors and find the vector}
\\\text{cross product of the 2 vectors}
\\(3,1,-2)-(1,-2,4)=(2,3,-6)
\\(2,-1,6)-(3,1,-2) = (-1,-2,8)
\\\therefore \overrightarrow{n} = \begin{vmatrix} i&j&k\\
2&3&-6\\
-1 & -2&8
\end{vmatrix}
\\\implies \overrightarrow{n}=12i-10j-k
\\\text{Next, we substitute the values of $\overrightarrow{n}$, $\overrightarrow{a}$ and $\overrightarrow{r}$ in equation(1), we have}
\\ 12x-10y-z=28"